OpenStudy (anonymous):

Convert the following percent to a decimal. 80%

OpenStudy (utterly_confuzzled):

80% = 80/100 = 8/10 = 0.8

OpenStudy (carlyleukhardt):

80% as a decimal is .8

OpenStudy (anonymous):

ok thank u both

OpenStudy (carlyleukhardt):

no prob

OpenStudy (texaschic101):

when turning a percent to a decimal...start at the percent, and move it 2 spaces to the left and change to a decimal.

OpenStudy (anonymous):

ok :)

OpenStudy (triciaal):

percent means per 100 80/100 =

OpenStudy (anonymous):

What is the range of the following set of data? 54, 43, 56, 51, 49, 48, 46, 59, 46

OpenStudy (texaschic101):

let me rephrase that... because if you had 32.2%......start at the decimal, and go 2 spaces to the left....so this example would be .322 range is the largest number minus the smallest number

OpenStudy (triciaal):

from the lowest number to the highest number

OpenStudy (anonymous):

idk how to find range

OpenStudy (texaschic101):

what is the highest number in that set ?

OpenStudy (texaschic101):

and what is the lowest number in that set ?

OpenStudy (texaschic101):

subtract the lowest from the highest and that is your range

OpenStudy (anonymous):

ok